Examine the following code for (int x = 1; x < 34; x = x + 2) { x = x % 40; x = x - 1; MessageBox.Show(x.ToString()); } How many times is this loop going to iterate (body is going to execute)
Answer #1

Ans : 33 Times

Explanation :

0 x = 1, x = 1%40 = 1, x = x-1 , x = 1-1 = 0
1 x = 2 x = 2%40 = 2, x = x-1 , x = 2-1 = 1
2 x = 3 x = 3%40 = 3, x = x-1 , x = 3-1 = 2
3 x = 4, x = 4%40 = 4, x = x-1 , x = 4-1 = 3
4 x = 5, x = 5%40 = 5, x = x-1 , x = 5-1 = 4
5 x = 6, x = 6%40 = 6, x = x-1 , x = 6-1 = 5
6 x = 7, x = 7%40 = 7, x = x-1 , x = 7-1 = 6
7 x = 8, x = 8%40 = 8, x = x-1 , x = 8-1 = 7
8 x = 9, x = 9%40 = 9, x = x-1 , x = 9-1 = 8
9 x = 10, x = 10%40 = 10, x = x-1 , x = 10-1 = 9
10 x = 11, x = 11%40 = 11, x = x-1 , x = 11-1 = 10
11 x = 12, x = 12%40 = 12, x = x-1 , x = 12-1 = 11
12 x = 13, x = 13%40 = 13, x = x-1 , x = 13-1 = 12
13 x = 14, x = 14%40 = 14, x = x-1 , x = 14-1 = 13
14 x = 15, x = 15%40 = 15, x = x-1 , x = 15-1 = 14
15 x = 16, x = 16%40 = 16, x = x-1 , x = 16-1 = 15
16 x = 17, x = 17%40 = 17, x = x-1 , x = 17-1 = 16
17 x = 18, x = 18%40 = 18, x = x-1 , x = 18-1 = 17
18 x = 19, x = 19%40 = 19, x = x-1 , x = 19-1 = 18
19 x = 20, x = 20%40 = 20, x = x-1 , x = 20-1 = 19
20 x = 21, x = 21%40 = 21, x = x-1 , x = 21-1 = 20
21 x = 22, x = 22%40 = 22, x = x-1 , x = 22-1 = 21
22 x = 23, x = 23%40 = 23, x = x-1 , x = 23-1 = 22
23 x = 24, x = 24%40 = 24, x = x-1 , x = 24-1 = 23
24 x = 25, x = 25%40 = 25, x = x-1 , x = 25-1 = 24
25 x = 26, x = 26%40 = 26, x = x-1 , x = 26-1 = 25
26 x = 27, x = 27%40 = 27, x = x-1 , x = 27-1 = 26
27 x = 28, x = 28%40 = 28, x = x-1 , x = 28-1 = 27
28 x = 29, x = 29%40 = 29, x = x-1 , x = 29-1 = 28
29 x = 30, x = 30%40 = 30, x = x-1 , x = 30-1 = 29
30 x = 31, x = 31%40 = 31, x = x-1 , x = 31-1 = 30
31 x = 32, x = 32%40 = 32, x = x-1 , x = 32-1 = 31
32 x = 33, x = 33%40 = 33, x = x-1 , x = 33-1 = 32

Do ask if any doubt. Please upvote.